# Robust Estimation of Polyserial Correlation

### v1

@@ D.1 Contamination through gross errors
Figure D.4: Visualizations of the results of the “gross error” simulation design in Section D.1 for teh considered estimators across 5,000 repetitions, at significance level $\gamma=0.05$.
@@
Figure D.5: For the “gross error” design: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge (top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom). The point estimators always converged, resulting in overplotted lines.
@@ D.2 Contamination through correlation shift
We generate $N=500$ observations from the described process and repeat this procedure 5,000 times. The same performance measures and estimators as in Section 7 are used. We stress that this simulation design has been used before by Welz et al., 2025b in the context of the polychoric correlation model.
@@
Figure D.8: For the “correlation shift” design: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge (top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom). The point estimators always converged, resulting in overplotted lines.
@@
Overall, this simulation study demonstrates that the magnitude of our proposed estimator’s robustness gain depends on the overlap between the contamination distribution $H_{X\eta}$ and true normal distribution ${P}_{X\eta}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$. If these distributions substantially overlap—like in correlation-shifted contamination with weak to moderate true correlation—then the robust estimator may not be able to clearly distinguish between contamination and regular observations since it makes no assumptions on the former, and is subsequently unable to provide a tangible robustness gain. On the other hand, when $H_{X\eta}$ and ${P}_{X\eta}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$ are sufficiently distinct—like in correlation-shifted contamination with strong true correlation or mean-shifted correlation (cf. Sections 7 and D.1)—then the robust estimator can identify the contaminated observations and subsequently downweigh them to achieve considerable robustness gains. In either scenario, we observe that the robust estimator is always at least as accurate as the ML estimator, and therefore constitutes an overall improvement in terms of robustness. Similar findings for a different robust estimator in the context of estimating polychoric correlation have been reported by Welz et al., 2025b .
@@ D.3.1 Preliminaries
Although conceptually different, robustness to partial and distributional misspecification frameworks are “practically synonymous notions” (Huber & Ronchetti,, 2009, p. 4). Thus, despite our robust estimator being designed for a fundamentally different type of model misspecification—partial rather than distributional misspecification—it is a relevant question whether or not our estimator can also offer enhanced robustness to distributional misspecification. Welz et al., 2025b (, p. 32) argue that the potential for robustness gains of partial-misspecification-robust estimators in distributional misspecification depends on the properties of the unknown sampling distribution $G$. If the nonnormal $G$ can be decently approximated by mixture between a normal distribution and some other distribution—thereby approximating the partial misspecification framework in (11)—then our robust estimator should perform reasonably well. If $G$ cannot be approximated by such a mixture, neither ML nor our estimator may be expected to perform well.
@@ D.3.2 Simulation study
Figure D.11: For the “distributional misspecification” design: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge (top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom).
@@
Figure D.10 and Table D.1 summarize the simulation results. When the joint distribution $G$ is a Clayton copula, the MLE exhibits a bias of about 0.066-0.066, whereas all robust estimators are less biased (in absolute magnitude). The MLE’s coverage is about 60%, but it also has the widest confidence intervals (wider then the robust estimators). In contrast, despite having narrower confidence intervals, the estimators with $\alpha>0.1$ achieve a coverage of about 60% or higher. Especially the choices $\alpha=0.5,0.75$ are remarkably accurate with coverages of about 95% and 88%, respectively. Conversely, when distribution $G$ is a Gumbel copula, all estimators exhibit a notable bias of similar magnitude. The slightly higher coverage of the robust estimators is primarily due to wider confidence intervals. Thus, in this setting, the robust estimators do not improve upon the MLE.
@@
We noticed that in a small number of repetitions, the robust estimator did not numerically converge. This situation occurs when multiple response options of $Y$ are almost never chosen (e.g., only once or twice in the sample). In such a situation, the robust estimator may attempt to eliminate a threshold by either pushing the first or last threshold to $\pm\infty$ or pulling adjacent thresholds together as closely as numerically possible. Similar behavior has been observed for a robust estimator of the polychoric correlation model (Welz et al., 2025b, ). We follow Welz et al., 2025b by identifying numerical instability through adjacent thresholds being unreasonably far from each other, namely by a minimum distance of 3.92. Under the polyserial model, this distance covers as much as 95% of all probability mass of the standard normal marginal distribution of $\eta$. We subsequently omitted the estimates where such nonconvergence occurs from the analyses in this section. Figure D.11 summarizes the numerical nonconvergence statistics. The highest fraction of nonconvergent cases occurs for the choice $\alpha=0.5$ with about 3.5%, which is a reasonably small value considering that our proposed estimator is not designed for distributional misspecification.
@@
Overall, the results suggest that for population correlation 0.7, the Clayton copula might be reasonably well-approximable by a mixture between a normal distribution and some other distribution, but not the Gumbel copula. Figure D.9 indicates that while the two copulas are of comparable shape, the Clayton copula has substantially more probability mass in its center than the Gumbel distribution, which may enable it to be somewhat approximable by a mixture involving a normal distribution. A similar result for the Clayton copula has been found in Welz et al., 2025b for distributional misspecification of the polychoric correlation model.

### v2

@@ D.1 Contamination through gross errors
Figure D.5: Visualizations of the results of the “gross error” simulation design in Section D.1 for the considered estimators across 5,000 repetitions, at significance level $\gamma=0.05$.
@@
Figure D.6: For the “gross error” design: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge—neither with the unconstrained BFGS nor the constrained Nelder-Mead algorithm—(top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom). The point estimators always converged, resulting in overplotted lines. For $\varepsilon>0$, no standard errors could ever be computed for ML, reflecting the instability of ML estimation in the presence of gross errors.
@@ D.2 Contamination through correlation shift
We generate $N=500$ observations from the described process and repeat this procedure 5,000 times. The same performance measures and estimators as in Section 7 are used. We stress that this simulation design has been used before by Welz et al., (2026) in the context of the polychoric correlation model.
@@
Figure D.9: For the “correlation shift” design: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge—neither with the unconstrained BFGS nor the constrained Nelder-Mead algorithm—(top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom). The point estimators always converged, resulting in overplotted lines.
@@
Overall, this simulation study demonstrates that the magnitude of our proposed estimator’s robustness gain depends on the overlap between the contamination distribution $H_{X\eta}$ and true normal distribution ${P}_{X\eta}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$. If these distributions substantially overlap—like in correlation-shifted contamination with weak to moderate true correlation—then the robust estimator may not be able to clearly distinguish between contamination and regular observations since it makes no assumptions on the former, and is subsequently unable to provide a tangible robustness gain. On the other hand, when $H_{X\eta}$ and ${P}_{X\eta}\left(\cdot,\cdot;\bm{\theta}_{*}\right)$ are sufficiently distinct—like in correlation-shifted contamination with strong true correlation or mean-shifted correlation (cf. Sections 7 and D.1)—then the robust estimator can identify the contaminated observations and subsequently downweigh them to achieve considerable robustness gains. In either scenario, we observe that the robust estimator is always at least as accurate as the ML estimator, and therefore constitutes an overall improvement in terms of robustness. Similar findings for a different robust estimator in the context of estimating polychoric correlation have been reported by Welz et al., (2026).
@@ D.3.1 Preliminaries
Although conceptually different, robustness to partial and distributional misspecification frameworks are “practically synonymous notions” (Huber & Ronchetti,, 2009, p. 4). Thus, despite our robust estimator being designed for a fundamentally different type of model misspecification—partial rather than distributional misspecification—it is a relevant question whether or not our estimator can also offer enhanced robustness to distributional misspecification. Welz et al., (2026, p. 32) argue that the potential for robustness gains of partial-misspecification-robust estimators in distributional misspecification depends on the properties of the unknown sampling distribution $G$. If the nonnormal $G$ can be decently approximated by mixture between a normal distribution and some other distribution—thereby approximating the partial misspecification framework in (8)—then our robust estimator should perform reasonably well. If $G$ cannot be approximated by such a mixture, neither ML nor our estimator may be expected to perform well.
@@ D.3.2 Simulation study
Figure D.12: For the “distributional misspecification” design: Fraction of the 5,000 simulation repetitions in which a point estimator did not numerically converge—neither with the unconstrained BFGS nor the constrained Nelder-Mead algorithm—(top), and its associated standard error could not be computed due to singularity of $\widehat{\bm{J}}_{N}\left(\widehat{\bm{\theta}}_{N}\right)$ in the covariance matrix estimator (bottom).
@@
Figure D.11 and Table D.1 summarize the simulation results. When the joint distribution $G$ is a Clayton copula, the MLE exhibits a bias of about 0.066-0.066, whereas all robust estimators are less biased (in absolute magnitude). The MLE’s coverage is about 60%, but it also has the widest confidence intervals (wider then the robust estimators). In contrast, despite having narrower confidence intervals, the estimators with $\alpha>0.1$ achieve a coverage of about 60% or higher. Especially the choices $\alpha=0.5,0.75$ are remarkably accurate with coverages of about 95% and 87%, respectively. Conversely, when distribution $G$ is a Gumbel copula, all estimators exhibit a notable bias of similar magnitude. The slightly higher coverage of the robust estimators is primarily due to wider confidence intervals. Thus, in this setting, the robust estimators do not improve upon the MLE.
@@
We noticed that in a small number of repetitions, the robust estimator did not numerically converge. This situation occurs when multiple response options of $Y$ are almost never chosen (e.g., only once or twice in the sample). In such a situation, the robust estimator may attempt to eliminate a threshold by either pushing the first or last threshold to $\pm\infty$ or pulling adjacent thresholds together as closely as numerically possible. Similar behavior has been observed for a robust estimator of the polychoric correlation model (Welz et al.,, 2026). We follow Welz et al., (2026) by identifying numerical instability through adjacent thresholds being unreasonably far from each other, namely by a minimum distance of 3.92. Under the polyserial model, this distance covers as much as 95% of all probability mass of the standard normal marginal distribution of $\eta$. We subsequently omitted the estimates where such nonconvergence occurs from the analyses in this section. Figure D.12 summarizes the numerical nonconvergence statistics. The highest fraction of nonconvergent cases occurs for the choice $\alpha=0.{\color[rgb]{0,0,0}7}5$ with about 4%, which is a reasonably small value considering that our proposed estimator is not designed for distributional misspecification.
@@
Overall, the results suggest that for population correlation 0.7, the Clayton copula might be reasonably well-approximable by a mixture between a normal distribution and some other distribution, but not the Gumbel copula. Figure D.10 indicates that while the two copulas are of comparable shape, the Clayton copula has substantially more probability mass in its center than the Gumbel distribution, which may enable it to be somewhat approximable by a mixture involving a normal distribution. A similar result for the Clayton copula has been found in Welz et al., (2026) for distributional misspecification of the polychoric correlation model.
@@
As a practical consequence, we recommend applied researchers to stay wary of potential partially-latent nonnormality even when ML and our robust estimator produce very similar point estimates. While such similarity may be due to normality holding true (in which case both estimators yield asymptotically equivalent estimates), it may also be due to distributional misspecification from a specific nonnormal distribution for which the robust estimator cannot improve over ML (such as the Gumbel copula considered in this simulation).iiiA third possibility is that partially-latent normality is violated in the $\eta$-dimension, but the misspecification is inconsequential because the ensuing density of the observed $(X,Y)$ is equal to the polyserial model distribution in (4) (see Section 4). In this case, both ML and robust estimators will yield accurate point estimates despite the model being technically misspecified. Hence, in this theoretical case, there are no adverse effects for point estimation with either estimator. We refer to Grønneberg et al., (2020) and Foldnes & Grønneberg, (2020) for detailed discussions on testing for underlying normality from discretized variables. To investigate the possibility of distributional misspecification whenever ML and robust estimator yield similar estimates, a $\chi^{2}$-test for partially-latent normality may be considered, which is, for example, implemented in the R package polycor (Fox,, 2022).iiiiiiRecently, Foldnes & Grønneberg, (2020) have proposed a bootstrap test for testing underlying multivariate normality when all observed variables are ordinal. Therefore, the bootstrap test as described in Foldnes & Grønneberg, (2020) is not directly applicable to mixed data. However, we believe that it should be possible to extend their methodology to mixed data. We leave this to future research.